# ”It’s Mentally Painful to Stop”: Design Opportunities in In-Situ Self-Management Technology for People with Obsessive-Compulsive Disorder

### v1

@@ Abstract.
Obsessive-compulsive disorder (OCD) is a mental health condition significantly affecting people’s quality of life. Although OCD can be effectively treated with evidence-based therapy (e.g., exposure and response prevention), Managing OCD symptoms independently, as an indispensable part of successful treatment, remains challenging due to fear confrontation and lack of appropriate support. We aim to comprehensively understand the challenges and needs in OCD self-management from the perspectives of both people with OCD and OCD therapists. Through interviews with 10 participants with diverse OCD conditions and seven therapists, we characterized different OCD symptoms, typical triggering factors, strategies, technology use, and barriers both inside and outside of therapy. Our findings highlighted gaps between OCD self-management needs and currently available support. Building on these insights, we suggest in-situ self-management technologies for OCD, including personalized symptom tracking, in-situ interventions, and discuss how OCD-specific privacy and social needs can be fulfilled with technology and beyond.
mental health, OCD, Obsessive-Compulsive Disorder, Exposure and Response Prevention, Acceptance and Commitment Therapy, in-situ intervention
copyright: noneccs: Human-centered computing Empirical studies in accessibilityccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Obsessive-Compulsive Disorder (OCD) is a mental health disorder affecting approximately 2% of world population (Sasson et al., 1997). People with OCD experience uncontrollable, recurring thoughts (obsessions) usually triggered by specific events (triggers), and engage in repetitive behaviors (compulsions) to relief the anxiety caused by obsessions (of Mental Health, NIMH). Contrary to the common misconception that OCD is simply ’being clean’ or ’focused on order’ (Spencer and Carel, 2021), OCD symptoms are highly diverse (Hirschtritt et al., 2017) and can significantly affect quality of life especially when compulsions become severe enough to disrupt daily activities (Macy et al., 2013; Eisen et al., 2006; Subramaniam et al., 2013).
@@
Coping with OCD symptoms independently in daily life is crucial but challenging as real-life triggers can be too overwhelming to confront during intense OCD episodes without suitable guidance (Lind et al., 2013; Hezel and Simpson, 2019; Wheaton and Chen, 2021; Ojalehto et al., 2020). Many individuals with OCD seek professional support through cognitive behavioral therapy (CBT), a type of psychotherapy that help people identify and change unhelpful thoughts and behavior patterns (Beck, 2011). Despite its effectiveness (Yule and Whittal, 2017; Blakey et al., 2017; Hirschtritt et al., 2017), over 50% of people with OCD lack access to the therapy treatment (García-Soriano et al., 2014; Ruscio et al., 2010; Chong et al., 2012; Mayerovitch et al., 2003). Moreover, many who complete the treatment continue to experience residual symptoms and struggle to apply the skills learned in therapy to their daily lives (Eddy et al., 2004). These challenges highlight the needs for effective self-management technology that supports people with OCD in-situ beyond traditional therapy.
@@
Recent advances in AI and sensing techniques have shown great potential for in-situ self-management for OCD—through intelligent scene understanding and behavior recognition (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024; Liu et al., 2024), such technology can detect triggering events and compulsive behaviors, enabling timely, in-situ interventions to help people with OCD address obsessions and resist compulsions during OCD flare-ups (Burchard et al., 2022; Olbrich et al., 2016). Despite the potential, research on in-situ intervention technologies for OCD remains limited and primarily focuses on specific symptoms, including recognizing excessive hand-washing on a smartwatch (Burchard et al., 2022) and detecting excessive checking behaviors via a short moving distance over an extended period of time based on GPS location (Olbrich et al., 2016). Given the complex and heterogeneous nature of OCD (Hirschtritt et al., 2017), timely research is needed to deeply understand the diverse personal experiences of people with OCD, uncover their challenges and needs, and identify suitable intervention feedback, thus inspiring effective, generalizable technology design.
To fill this gap, we contribute the first comprehensive study with both people with OCD and OCD professionals. We first interviewed 10 participants with diverse OCD symptoms about their detailed experiences of OCD, symptom self-management strategies, persisting challenges, and desired technology to assist with self-management. To complement these personal insights, we further conducted an interview with seven mental health care providers who specialized in OCD to learn about their treatment strategies throughout different stages of OCD therapy and technology recommendations to facilitate symptom management outside of therapy context, guiding the feedback design for in-situ intervention technology for OCD self-management.
Our studies detailed OCD people’s individualized triggers and compulsions and identified typical properties that predominantly affect trigger intensity. We found consistency between self-management strategies used by people with OCD and those taught by OCD therapists, while also uncovering nuances in treatment strategies and use of technology. By combining insights from both studies, we highlight gaps in OCD self-management despite currently available resources. By triangulating the findings from both studies, we derive technology design implications from trigger and compulsion detection, in-situ feedback design, and privacy perspectives. Beyond in-situ self-management, some implications can also benefit OCD therapy, such as supporting out-of-session practice.
In summary, our research contributes: (1) the first study that thoroughly understand in-situ self-management strategies and needs from perspectives of both people with OCD and OCD professionals (2) rich qualitative data that reveals detailed personal experiences of people with OCD to inspire in-situ technology design (3) design implications that focused on input (OCD trigger and compulsion recognition) and output (feedback design) of in-situ intervention technology and OCD-specific privacy and social needs.
@@ 2. Related Work
2.1. Obsessive-Compulsive Disorder & Therapy Treatments
2.1.1. What is OCD
@@
OCD is highly heterogeneous. Yale-Brown Obsessive Compulsive Scale (Y-BOCS) Symptom Checklist (Goodman et al., 1989) organizes different OCD symptoms into eight obsession categories and seven compulsion categories (Goodman et al., 1989; McKay et al., 2004). Common obsession themes include fear of contamination, persistent doubting, violent or sexual intrusive thoughts. In response to these obsessions, the compulsive behaviors can be excessive hand washing, repeated checking (e.g., whether the oven is off), and repeated neutralizing thoughts (e.g., “I’m not a bad person”), respectively (Hirschtritt et al., 2017). Unlike obsessions and compulsions, the diversity of triggers has been overlooked in OCD research. Existing studies focus on generic triggers for specific symptoms (Mataix-Cols et al., 2009) or describe OCD episodes without clearly distinguishing triggers (Knapton, 2016). Since triggers are highly personalized and important in characterizing OCD symptoms (Baioui et al., 2013), it is crucial to further understand and categorize triggers by exploring OCD people’s personal experiences.
2.1.2. Common Therapy Treatments for OCD
To combat OCD, cognitive behavioral therapy (CBT) is the most effective method (Hirschtritt et al., 2017), focusing on identifying and changing patients’ unwanted thinking and behavioral patterns (Beck, 2011). CBT includes different forms, such as exposure and response prevention (ERP) therapy, cognitive therapy (CT), and acceptance and commitment therapy (ACT) (Björgvinsson et al., 2007).
ERP therapy has been widely used to facilitate behavioral changes (Blakey et al., 2017; Björgvinsson et al., 2007; Hirschtritt et al., 2017). It is a psychotherapy that gradually exposes individuals with OCD to triggers that provoke obsessions while simultaneously discouraging compulsions (Blakey et al., 2017; Hirschtritt et al., 2017). To decide the exposure tasks, an exposure hierarchy is built to cover the clients’ OCD triggers ranked based on the intensity of distress they cause. In between sessions, clients are also expected to complete exposure practices to achieve better outcome. Once the client completes the most difficult exposure successfully, the treatment can be concluded. However, ERP’s nature of ‘facing the fears’ can distress the clients and be time consuming (Patel et al., 2021), hence resulting in high dropout rate (Ong et al., 2016).
To facilitate cognitive change, cognitive therapy (CT) is conducted to correct unhelpful and dysfunctional beliefs that lead to obsessions. Strategies are commonly developed from the underlying characteristics of the dysfunctional beliefs, such as overestimation of threat. An example strategy can be showing the client that the catastrophic outcome is a consequence of many steps so its probability is much lower than they expect (Yule and Whittal, 2017). Prior research shows that CT can achieve comparable effectiveness to ERP (Yule and Whittal, 2017), and that CT is better tolerated than ERP with lower dropout rate (Yule and Whittal, 2017; Björgvinsson et al., 2007).
More recently, acceptance and commitment therapy (ACT), as a newly developed form of CBT, has been applied to the treatment of OCD and can significantly reduce OCD symptoms (Philip and Cherian, 2021; Evey and Steinman, 2023). Unlike traditional CBT (e.g., ERP), ACT encourages people to accept the presence of obsessions and the distress resulting from them, without attempting to change these feelings through compulsive behaviors (Smith et al., 2017). It aims to improve the ability of people with OCD to stay at the present, not attending to previous or future experiences (Smith et al., 2017; Philip and Cherian, 2021; Evey and Steinman, 2023). In ACT, people with OCD are guided to identify the values that are more important to them and encouraged to commit to behaviors that are consistent to these values. For example, a person with OCD who has identified spending time with family as a value may find excessive hand washing interfering this important event of life. Research has shown that combining ERP and ACT can further improve people’s ability to respond to more intense obsessions (Ong et al., 2020).
In summary, CBT has proven to be effective in addressing OCD symptoms from both obsession and compulsion aspects. However, it is not available 24/7 whenever people need support to manage their issues (Blakey et al., 2017). The homework between sessions (e.g., symptom self-monitoring, self-guided exposure practices)—as an important part of CBT—could be difficult for clients to comply, affecting the outcome of the therapy (Wheaton and Chen, 2021).
@@ 2.2. Technology for OCD Symptom Self-Management
To overcome the challenges of OCD symptom self-management, a variety of different forms of new technology have been developed and evaluated. Existing efforts have been focused on symptom and behavior monitoring(Khan et al., 2011; Frank et al., 2023), proactive skill building(Khan et al., 2011; Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010), and in-situ intervention (Burchard et al., 2022; Olbrich et al., 2016). We discuss prior work on self-symptom management following this taxonomy.
@@ 2.2.1. Symptom monitoring
Symptom monitoring is an effective approach to supporting self-management for various mental health conditions, such as anxiety disorder and depression (Kruzan et al., 2023), due to its potential of improving a person’s symptom (e.g., thoughts, behavior) awareness (Kruzan et al., 2023). In the context of OCD, prior work has explored two ways of self-monitoring, (1) self-report, and (2) passive sensing.
Ecological Momentary Assessment (EMA) — a research approach in clinical psychology to periodically prompting for people’s behaviors and experiences in their natural environments — has been broadly used to track and assess an OCD person’s symptoms (Frank et al., 2023; Brown et al., 2020; Tilley and Rees, 2014; Rupp et al., 2020), and demonstrated high acceptability, practicability, and representativeness (Frank et al., 2023). However, research has shown that overly relying on self-report approach can result in noisy (Frank et al., 2023) and missing data.
In contrast, passive sensing enables rich data collection without burdening users. Researchers have investigated passive symptom monitoring using various sensors (Thierfelder et al., 2022; Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022). For example, Thierfelder et al. (Thierfelder et al., 2022) monitored OCD-induced stress by measuring heart rate, hand activity and gaze fixation, and demonstrated that stress and anxiety detection is possible using multimodal sensors. Other studies measured sleep quality of people with OCD using actigraphy (Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022). While symptom monitoring can improve symptom awareness, prior work has primarily focused on monitoring users’ internal states instead of contextual information which is important in characterizing OCD symptoms (Llorens-Aguilar et al., 2021). In addition, without proper knowledge and coping skills for symptom management, as well as actionable insights derived from collected data, the effectiveness of symptom monitoring can be further limited.
@@ 2.2.2. Proactive Skill Building
To support out-of-session symptom management skill building, smartphone-based intervention has been developed that incorporate evidence-based therapy treatment in peoples’ everyday life (Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010). LiveOCDFree is an app that guides the user to set up and conduct self-administered ERP by creating an exposure hierarchy, and identifying specific exposures to practice (e.g. touching doorknobs). The user is able to set reminders for ERP practice and monitor their progress. An evaluation with 21 participants demonstrated significant improvement in OCD and anxiety symptoms after using the app (Boisseau et al., 2017). Similarly, NOCD supported self-administered ERP between therapy sessions incorporating ACT strategies. A study with 2,069 participants showed that NOCD effectively reduced OCD symptoms (Feusner et al., 2021). However, there are limitations with self-administered ERP such as inflexible template for specifying obsessions and compulsions which restricts user autonomy (Hong et al., 2020). Prior work has also reported clinical significance of web-based CBT in reducing OCD symptoms (Ferreri et al., 2019), such as BT Steps, a program that guides users with OCD to learn symptom management skills (Schure et al., 2020).
These technologies effectively supported OCD people’s symptom management outside of therapy. However, they all require OCD clients to proactively practice to build up coping skills, instead of providing targeted support for clients to respond to intense real life triggers, which can be difficult to manage without therapists’ help (Lind et al., 2013). In contrast, in-situ intervention technologies that implement therapeutic strategies (e.g., exposure practice) along with real life events (Kim et al., 2024) can potentially assist with OCD treatment in a more effective manner.
2.2.3. In-Situ Intervention
As opposed to pre-scheduled therapy sessions, the increasingly powerful mobile and sensing technologies allow for more flexible and intelligent mental well-being interventions. Such systems can recognize and adapt to users’ internal mental states and physical context in daily life to provide just-in-time mental support — an intervention design that provides the right type and amount of support, at the right time, in the context that the person needs it most and is most likely to be receptive(Nahum-Shani et al., 2018).
Such interventions have been applied to treat different mental health conditions, such as substance use disorder (Gustafson et al., 2014), depression (Burns et al., 2011), and schizophrenia (Ben-Zeev et al., 2014, 2013). For example, ACHESS (Gustafson et al., 2014) is a mobile application that tracks the GPS location of a user with alcohol use disorders and provides text-based support when they go to high-risk locations (e.g., a bar). An evaluation with 349 participants shows that ACHESS significantly reduced risky drinking days.
There has been limited exploration on in-situ intervention for OCD. Burchard et al. (Burchard et al., 2022) develops a deep-learning-based system to detect excessive hand-washing—a typical compulsion in contamination OCD—using the inertial motion sensor data on a smartwatch. However, although the system supports real-time symptom recognition, it does not provide in-situ feedback to the user. Olbrich et al. (Olbrich et al., 2016) develops a smartphone application Geo-Feedback that actively detects the user’s excessive checking behavior indicated by short moving distance over an extended period of time via GPS location and evokes a selected alarm when the excessive checking behavior is detected. A case study with a person with severe OCD indicates the effectiveness of the application. These examples have demonstrated initial success in designing real-time interventions for people with OCD. However, they are purely focused on physical compulsions and highly limited in context, hindering generalization to other OCD conditions. More research is needed to guide the design of in-situ intervention technology, including both the detection of OCD triggers and the design of effective and acceptable interventions consistent with therapeutic advice.
3. Study 1: Interviewing People with OCD
We first conducted semi-structured interviews with 10 participants with OCD. We aimed to comprehensively understand the impact of OCD (e.g., obsessions, compulsions, triggers) on their daily life, with an ultimate goal of distilling the needs and challenges of OCD symptom self-management and informing the design of in-situ intervention. The study was approved by Institutional Review Board (IRB) at our university.
3.1. Method
3.1.1. Participants
We recruited 10 participants (7 female and 3 male) whose ages ranged from 19 to 47 ($Mean$ = 28.5, $SD$ = 8.5). Participants’ demographic information and OCD diagnosis are detailed in Table 1. We recruited participants from online OCD support groups and the research email service at our university. Participants were screened via an email to (1) inform them of the interview content to ensure they would be comfortable participating in this study, and (2) check their eligibility. A participant was eligible if they were (1) at least 18 years old, and (2) diagnosed with mild to moderate OCD and were willing to disclose their OCD symptoms with us. We did not recruit people experiencing severe OCD at the time of the study, because our interview questions can be triggering without adequate ability to manage symptoms independently. Participants were compensated for \$20 per hour. At the time of the study, all participants except P6 (from India) were located in the United States.

### v2

# “It was Mentally Painful to Try and Stop”: Design Opportunities for Just-in-Time Interventions for People with Obsessive-Compulsive Disorder in the Real World

@@ Abstract.
Obsessive-compulsive disorder (OCD) is a mental health condition that significantly impacts people’s quality of life. While evidence-based therapies such as exposure and response prevention (ERP) can be effective, managing OCD symptoms in everyday life—an essential part of treatment and independent living—remains challenging due to fear confrontation and lack of appropriate support. To better understand the challenges and needs in OCD self-management, we conducted interviews with 10 participants with diverse OCD conditions and seven therapists specializing in OCD treatment. Through these interviews, we explored the characteristics of participants’ triggers and how they shaped their compulsions, and uncovered key coping strategies across different stages of OCD episodes. Our findings highlight critical gaps between OCD self-management needs and currently available support. Building on these insights, we propose design opportunities for just-in-time self-management technologies for OCD, including personalized symptom tracking, just-in-time interventions, and support for OCD-specific privacy and social needs—through technology and beyond.
mental health, OCD, Obsessive-Compulsive Disorder, Exposure and Response Prevention, Acceptance and Commitment Therapy, just-in-time intervention
journalyear: 2025copyright: ccconference: The 27th International ACM SIGACCESS Conference on Computers and Accessibility; October 26–29, 2025; Denver, CO, USAbooktitle: The 27th International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’25), October 26–29, 2025, Denver, CO, USAdoi: 10.1145/3663547.3746394isbn: 979-8-4007-0676-9/2025/10ccs: Human-centered computing Empirical studies in accessibilityccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Obsessive-Compulsive Disorder (OCD) is a mental health disorder affecting approximately 2% of world population (Sasson et al., 1997). People with OCD experience uncontrollable, recurring thoughts (obsessions) usually triggered by specific events (triggers), and engage in repetitive behaviors (compulsions) to relieve the anxiety caused by obsessions (of Mental Health, NIMH). Contrary to the common misconception that OCD is simply ‘being clean’ or ‘focused on order’ (Spencer and Carel, 2021), OCD symptoms are highly diverse (Hirschtritt et al., 2017) and can significantly affect quality of life especially when compulsions become severe enough to disrupt daily activities (Macy et al., 2013; Eisen et al., 2006; Subramaniam et al., 2013).
@@
Coping with OCD symptoms independently in daily life is crucial but challenging as real-life triggers can be too overwhelming to confront without suitable guidance (Lind et al., 2013; Hezel and Simpson, 2019; Wheaton and Chen, 2021; Ojalehto et al., 2020). Many people with OCD seek professional support through cognitive behavioral therapy (CBT) to receive help in identifying and changing obsessive thoughts and behavior patterns (Beck, 2011). Despite its effectiveness (Yule and Whittal, 2017; Blakey et al., 2017; Hirschtritt et al., 2017), over 50% of people with OCD lack access to the therapy treatment (Garc´ıa-Soriano et al., 2014; Ruscio et al., 2010; Chong et al., 2012; Mayerovitch et al., 2003). Moreover, many who complete the treatment continue to experience residual symptoms and struggle to apply the skills learned in therapy to their daily lives (Eddy et al., 2004). These challenges highlight the needs for effective just-in-time self-management technology that supports people with OCD in real life beyond traditional therapy.
@@
Recent advances in AI and sensing techniques have shown great potential for just-in-time OCD interventions. Through intelligent scene understanding and behavior recognition (Xu et al., 2024; Xie et al., 2024; Bordes et al., 2024; Liu et al., 2024), such technology can potentially detect triggering events and compulsive behaviors and generate timely, in-situ interventions to help people with OCD address obsessions and resist compulsions during OCD flare-ups (Burchard et al., 2022; Olbrich et al., 2016). Despite the potential, HCI research on just-in-time intervention for OCD remains limited and primarily focuses on specific symptoms, for example, recognizing excessive checking behaviors when a short moving distance over an extended period of time is detected based on GPS location and providing audio alarm to remind the users (Olbrich et al., 2016). However, OCD is way more complex and heterogeneous with diverse triggers and symptoms (Hirschtritt et al., 2017). Timely research is needed to deeply understand the diverse triggering experiences of people with OCD, uncover their challenges and needs, and identify suitable intervention feedback, thus inspiring effective, generalizable technology design.
To fill this gap, we contribute the first interview study with ten people with OCD and seven OCD professionals to understand the diverse triggers and experiences of OCD, the persisting challenges faced by OCD individuals, real-life coping strategies, and desired (or recommended) technology for just-in-time interventions. Our study cover the perspectives of both stakeholders to identify the agreement and tension between people with OCD and the professionals, thus deriving most suitable intervention design guidelines that not only fulfill user needs but also align with professional practices.
Our study revealed unique findings towards OCD users’ experiences and preferences, as well as their tension with the professionals. Specifically, we found that instead of the trigger type in itself (e.g., a trash can), it was the different properties of triggers (e.g., fullness of the trashcan, whether there is a lid) that predominantly affect user experience. Moreover, besides certain consistency, we identified conflicts between the coping strategies adopted by people with OCD and those taught by OCD therapists. For example, the relaxing techniques (e.g., breathing practice) commonly used by OCD users in an OCD episode were not always recommended by the professionals. Our research highlights the gaps in OCD self-management despite currently available resources and derives technology design implications from trigger and compulsion detection, intervention feedback design, and ethics and privacy perspectives. Beyond just-in-time interventions in real life, some of our implications can also benefit OCD therapy, such as supporting out-of-session practice.
@@ 2. Related Work
2.1. Background of OCD
@@
OCD is highly heterogeneous. Yale-Brown Obsessive Compulsive Scale (Y-BOCS) Symptom Checklist (Goodman et al., 1989) organizes different OCD symptoms into eight obsession categories and seven compulsion categories (Goodman et al., 1989; McKay et al., 2004). Common obsession themes include fear of contamination, persistent doubting, violent or sexual intrusive thoughts. In response to these obsessions, the compulsive behaviors can be excessive hand washing, repeated checking (e.g., whether the oven is off), and repeated neutralizing thoughts (e.g., “I’m not a bad person”), respectively (Hirschtritt et al., 2017). However, unlike obsessions and compulsions, the diversity of triggers that cause obsessions and compulsions has been overlooked in OCD research. Existing studies focus on generic triggers for specific symptoms (Mataix-Cols et al., 2009) or describe OCD episodes without clearly distinguishing triggers (Knapton, 2016). Since triggers are highly personalized and important in characterizing OCD symptoms (Baioui et al., 2013), it is crucial to understand and categorize triggers by exploring OCD people’s personal experiences.
To combat OCD, cognitive behavioral therapy (CBT) is the most effective method (Hirschtritt et al., 2017), focusing on identifying and changing people’s unwanted thinking and behavioral patterns (Beck, 2011). CBT for OCD includes different forms, such as exposure and response prevention (ERP) therapy, cognitive therapy (CT), and acceptance and commitment therapy (ACT) (Björgvinsson et al., 2007). ERP helps people with OCD gradually face their OCD triggers while refraining from compulsions, aiming for behavioral change through repeated exposure to those triggers (Blakey et al., 2017; Björgvinsson et al., 2007; Hirschtritt et al., 2017). Cognitive Therapy (CT) focuses on identifying and restructuring dysfunctional beliefs that lead to obsessions, helping individuals reevaluate perceived catastrophic outcome (Yule and Whittal, 2017). Acceptance and Commitment Therapy (ACT), in contrast, encourages acceptance of obsessive thoughts without engaging in compulsions and promotes value-driven behavior to reduce the impact of OCD symptoms (Smith et al., 2017, 2017; Philip and Cherian, 2021; Evey and Steinman, 2023). CBT has proven to be effective in addressing OCD symptoms from both obsession and compulsion aspects. However, it is not available 24/7 whenever people need support to manage their issues (Blakey et al., 2017). The homework between sessions (e.g., symptom self-monitoring, self-guided exposure practices)—as an important part of CBT—could be difficult for clients to comply, affecting the outcome of therapy (Wheaton and Chen, 2021). Despite the drawbacks, the strategies used in these therapies can potentially bring insights for just-in-time interventions. Our research thus seize this opportunity to conduct studies with people with OCD to understand their experiences and coping strategies, triangulated with professionals’ perspectives, to inspire real-world intervention technologies beyond conventional therapy.
@@ 2.2. Technology for OCD Symptom Self-Management
To overcome the challenges caused by OCD in real life, different forms of technologies have been developed and evaluated. Existing efforts have mainly focused on symptom and behavior monitoring (Khan et al., 2011; Frank et al., 2023) and proactive skill building (Khan et al., 2011; Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010), but limited work has focused on just-in-time interventions for specific OCD symptoms (Burchard et al., 2022; Olbrich et al., 2016). We discuss prior work on OCD self-management technology below.
@@ 2.2.1. Symptom monitoring
Symptom monitoring is an effective approach to supporting mental health self-management, such as anxiety disorder and depression (Kruzan et al., 2023), due to its potential improvement in symptom (e.g., thoughts, behavior) awareness (Kruzan et al., 2023). Specifically in the context of OCD, prior work has explored two ways of self-monitoring, (1) self-report, and (2) passive sensing.
The most common self-report approach is via Ecological Momentary Assessment (EMA), which periodically prompts the users to track and assess their own symptoms in their natural environments (Frank et al., 2023; Brown et al., 2020; Tilley and Rees, 2014; Rupp et al., 2020). While this method has demonstrated high acceptability, practicability, and representativeness (Frank et al., 2023), research has shown that overly relying on self-report approach can result in noisy and missing data (Frank et al., 2023).
In contrast, passive sensing enables rich data collection without burdening users. Researchers have investigated passive symptom monitoring using various sensors (Thierfelder et al., 2022; Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022; Lønfeldt et al., 2023). For example, Thierfelder et al. (Thierfelder et al., 2022) monitored OCD-induced stress by measuring heart rate, hand activity and gaze fixation, and demonstrated the possibility of using multimodal sensory data to detect stress and anxiety. Other studies have focused on monitoring different aspects of OCD symptoms, including compulsion (e.g., excessive hand-washing) detection using inertial motion sensor data on a smartwatch (Wahl et al., 2022; Burchard et al., 2022), OCD episode tracking using blood volume pulse and skin temperature (Lønfeldt et al., 2023; Olesen et al., 2023), and sleep quality measurement using actigraphy (Alfano and Kim, 2011; Drummond et al., 2012; Coles et al., 2020; Cox and Olatunji, 2022). While symptom monitoring can improve symptom awareness, prior work has primarily focused on monitoring users’ internal states instead of contextual information (e.g., potential triggers in the environment), which is important in characterizing OCD symptoms (Llorens-Aguilar et al., 2021). In addition, without proper knowledge and coping skills for symptom management, the effectiveness of symptom monitoring can be further limited.
@@ 2.2.2. Proactive Skill Building
To support skill building outside therapy sessions, virtual therapy applications based on desktop or smartphone have been developed (Ferreri et al., 2019; Hong et al., 2020; Boisseau et al., 2017; Abramowitz et al., 2010). For example, LiveOCDFree is an app that guides the user to set up and conduct self-administered ERP therapy by creating an exposure hierarchy and identifying specific triggers to practice (e.g., touching doorknobs). The user is able to set reminders for ERP practice and monitor their progress. An evaluation with 21 participants demonstrated significant improvement in OCD and anxiety symptoms after using the app (Boisseau et al., 2017). Similarly, NOCD supported self-administered ERP between therapy sessions by incorporating ACT strategies. A study with 2,069 participants showed that it effectively reduced OCD symptoms (Feusner et al., 2021). Beyond conventional applications, recent advances in large language models (LLMs) further presents new opportunities in delivering personalized mental well-being support (Guo et al., 2024; Kian et al., 2024; Nie et al., 2024; Iftikhar et al., 2024). For example, Kian et al. (Kian et al., 2024) developed an LLM-powered assistive robot to deliver out-of-session CBT exercises for users with depression and achieved comparable effectiveness to traditional methods.
All above therapy simulation technologies require users to proactively practice to build up coping skills. However, triggers in real world could be much more unpredictable and intense, which can still be hard to manage even with certain skills established (Eddy et al., 2004). In contrast, targeted, just-in-time interventions that are adaptive to real world events (Kim et al., 2024) can potentially be more effective in real life.
2.2.3. Just-in-time Interventions
As opposed to pre-scheduled therapy sessions, the increasingly powerful mobile and sensing technologies allow for more flexible and intelligent mental well-being interventions. Such systems can recognize and adapt to users’ internal mental states and physical context in daily life and provide just-in-time support—an intervention design that delivers the right type and amount of support, adaptive to the timing and context when the user needs it. (Nahum-Shani et al., 2018). Just-in-time interventions have been applied to various mental health conditions, such as substance use disorder (Gustafson et al., 2014), depression (Burns et al., 2011; He et al., 2024; Zhao et al., 2023; Costa et al., 2016), schizophrenia (Ben-Zeev et al., 2014, 2013), and excessive phone use (Xu et al., 2022) For example, ACHESS (Gustafson et al., 2014) is a mobile app that tracks the GPS location of a user with alcohol use disorders and provides text-based reminder when they go to high-risk locations (e.g., a bar). An evaluation with 349 participants shows that ACHESS significantly reduced risky drinking days. Zhao et al. (Zhao et al., 2023) developed a wearable device that effectively reduced anxiety during high-stress situations through affective touch—a tactile intervention that renders a soft stroking sensation on the user’s forearm. Moreover, Xu et al. (Xu et al., 2022) designed TypeOut, which integrates value-based self-affirmation messages with a typing-based app unlock process, significantly reducing app usage and screen time.
Despite the exploration of just-in-time interventions for other mental health conditions, the work for OCD remains nascent. The only work we know is from Olbrich et al. (Olbrich et al., 2016), who develops a smartphone app Geo-Feedback that actively detects users’ excessive checking behaviors indicated by short moving distance over an extended period of time via GPS location and evokes an alarm when such excessive behaviors are detected. A case study with one user with severe OCD indicates the effectiveness of the app. While demonstrating the initial success of just-in-time interventions for people with OCD, this work focuses only on one specific OCD condition. More research is needed to understand the diverse, complex nature of OCD as well as user needs, thus deriving more generalizable guidelines for just-in-time intervention technology (including both trigger detection and intervention design) to support various OCD types in real life.
3. Method
To better understand how OCD (e.g., obsessions, compulsions, triggers) impacts people’s daily life and to identify effective just-in-time intervention strategies, we conducted in-depth interviews with 10 people with OCD (P1-P10) and seven OCD therapists (T1-T7). Our goal was to distill the triggering events, strategies, and challenges of OCD self-management in real life, ultimately informing the design of just-in-time intervention technologies.